OpenStudy (dancergirl45):

Given the function f(x) = 4(2)^x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other

OpenStudy (dancergirl45):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

Is it 4(2)x ????

OpenStudy (dancergirl45):

no it is hold on i will change it

OpenStudy (dancergirl45):

ok

OpenStudy (jango_in_dtown):

ok \[4.2^{x}\]

OpenStudy (jango_in_dtown):

right??

OpenStudy (jango_in_dtown):

average rate of change of f from x=a to x=b is \[\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (dancergirl45):

\[4(2)^{x}\]

OpenStudy (jango_in_dtown):

now for section A put b=2, a=1 in the above formula and tell me the answer

OpenStudy (dancergirl45):

wait what is f

OpenStudy (jango_in_dtown):

f(x)=given fuction

OpenStudy (dancergirl45):

so 1/1=1

OpenStudy (jango_in_dtown):

no why 1/1 ????

OpenStudy (dancergirl45):

well 2-1=1

OpenStudy (jango_in_dtown):

\[f(x)=4.2^{x}\]

OpenStudy (dancergirl45):

oh 7.4

OpenStudy (jango_in_dtown):



OpenStudy (dancergirl45):

4.2^ -- 1

OpenStudy (jango_in_dtown):

=8

OpenStudy (jango_in_dtown):

right??

OpenStudy (dancergirl45):

yes

OpenStudy (jango_in_dtown):

now for section B put b=4 and a=3 and tell me the answer

OpenStudy (dancergirl45):

3.2/1

OpenStudy (jango_in_dtown):

no